OpenStudy (amonoconnor):

When taking the anti-derivative of a definite integral, if the quantity of the integral/function = "36(2x + 1)^(-3)dx", does it become... 1) (36x)(2x + 1)^(-2) Or... 2) 36(2x + 1)^(-2) Or... 3) (36x)(2x + 1)^(-2)(2x^2 + x) ?? Any and all help is greatly appreciated! :)

ganeshie8 (ganeshie8):

try "u sub"

OpenStudy (amonoconnor):

... So you can't evaluate this by taking an anti derivative and plugging in the limits?

ganeshie8 (ganeshie8):

you can use advanced guessing

ganeshie8 (ganeshie8):

\(\large \int 36(2x + 1)^{-3}dx\) since it is of x^n form, you should get something like below : \(\large 36\frac{(2x + 1)^{-3+1}}{-3+1} \)

ganeshie8 (ganeshie8):

you need to fix it a bit, cuz its not just x in the parenthesis, its 2x+1 !

OpenStudy (amonoconnor):

Okay, so you don't add an exponent to the 36, and your don't use the chain rule (that would be backwards). Okay...

ganeshie8 (ganeshie8):

holdup - you still need to fix it, yes you need to use chain rule backwards

OpenStudy (amonoconnor):

Fix what? Which post?

ganeshie8 (ganeshie8):

\(\large \int 36(2x + 1)^{-3}dx\) since it is of x^n form, you should get something like below : \(\large 36\frac{(2x + 1)^{-3+1}}{-3+1} \) since its (2x+1) inside the parenthesis, divide above stuff by (2x+1)' = 2 \(\large 36\frac{(2x + 1)^{-3+1}}{2(-3+1)} + c \)

hartnn (hartnn):